# Impact of the carbon price on credit portfolio’s loss with stochastic collateral

### v1

@@ 5.1 Data
    Metropolitan France housing price index comes from OECD data and are from 1980 to 2021 (see OECD Stat, (2024) for details) in Base 2015. We renormalize in Base 2021.
@@ 5.2 Definition of the climate transition
We consider four deterministic transition scenarios giving four deterministic carbon price trajectories. The scenarios used come from the NGFS simulations, whose descriptions are given by NGFS, (2022) as follows:
@@
    Net Zero 2050 is an ambitious scenario that limits global warming to $1.5^{\circ}C$ through stringent climate policies and innovation, reaching net zero $\mathrm{CO}_{2}$ emissions around 2050. Some jurisdictions such as the US, EU and Japan reach net zero for all GHG by this point.
@@
For each scenario, we compute the carbon price $P_{carbon,0}$ in $t_{0}$ and the evolution rate $\eta_{\delta}$ as defined in (2.9).
@@
For the energy price, we consider electricity as the unique source of energy. Then, we assume a linear relation between the electricity and the carbon price inspired by Abrell et al., (2023), where a variation of the carbon price is linked withe the variation of the electricity by a the pass-through rate noted $k$. This means that $\mathfrak{f}^{elec}_{1}$ and $\mathfrak{f}^{elec}_{2}$ define in (3.32) are respectively $k$ and $P_{elec,0}-k\times P_{carbon,0}$. For France, we take the electricity price $P_{elec,0}=0.2161$ euro per Kilowatt-hour and $k=0.55$ (see Abrell et al., (2023)) ton per Kilowatt-hour. Its evolution is plotted in Figure 1(b).
@@ 5.3.2 Economic and housing pricing index (HPI) parameters
We keep the values of $\phi$, $\sigma$, $(\chi^{i})_{i\in\mathcal{I}}$, and $(\boldsymbol{\lambda}^{ji})_{i,j\in\mathcal{I}}$ already estimated. For the productivity process, we switched from a vector autoregressive model to an Ornstein-Uhlenbeck. We therefore calibrate ${\mu}$, ${\varsigma}$, ${\Sigma}$, and ${\Gamma}$ as detailed in Section 4.1.2 and we obtain the same results as in Sopgoui, (2024)[Section 4.3.1.].
@@ 5.4 Simulations and discussions
In the current simulations, since we are keeping the same data at the same frequency (annual), the main change is then to replace the VAR process by the O.-U. process. Therefore, the comments already made for (1) to (5) concerning the trends, the impact of the carbon price, the difference of scenarios, the relation between sectors, etc. do not change. We will focus here on the LGD and on the losses, with different type of collateral.
@@ 5.4.1 Impact of the carbon price on Loss Given Default
We compute here for $M=500$ simulations of the productivity processes $(\mathcal{Z},\mathcal{A})$, the loss given default of 4 loans with the same exposure but with 4 different financial assets collateral described in Table 3.
@@
We use the $M=500$ trajectories of the productivity processes $(\mathcal{Z},\mathcal{A})$ simulated above. We compute the loss given default of 4 loans with the same exposure but with the 4 buildings described in Table 5 as collateral.
@@
All the comments made for a financial asset as collateral are valid here: the presence of a collateral reduces LGD, that increases when the climate transition scenario becomes tougher.
@@ 5.4.2 Expected and unexpected loss
We can remark that, for each $k=0,\ldots,2$, firms $14k+$ 1, $24k+$ 2, $34k+$ 3, and $44k+$ 4 respectively belong to the Very High Emitting, High Emitting, Low Emitting, and Very Low Emitting groups. Moreover, we assume that
@@ Conclusion
Following Bouveret et al., (2023), we developed here a framework to quantify the impacts of the carbon price on a credit portfolio (expected and unexpected) losses, when the obligor companies as well as their guarantees belong to an economy subject to the climate transition declined by carbon price. We start by describing a closed economy, driven by a productivity following a multidimensional Ornstein-Uhlenbeck and subject to a climate transition modeled through a dynamic and deterministic carbon price, by a dynamic stochastic multisectoral. Then, by using the discounted cash flow methodology with the cash flows, following a stochastic differential equation, depending on the productivity as well as the carbon price, we evaluate the obligor value that helps us later on to compute its probability of over-indebtedness. We then turn to the bank’s loss in the event of a borrower’s over-indebtedness and if its loan is collateralized. When that is the case, the potential loss of the bank is written as the difference between the debt amount (EAD) and the collateral liquidated. We finally distinguish two types of collateral: either a financial asset or a building, both belonging to the economy so affected by the productivity and the carbon price. This work opens the door to many extensions as a finer modeling of the real estate market, taking into account other types of guarantees, modeling the unsecured loans that we assumed constant, modelling the impact of the carbon price on the exposure.

### v2

@@ 6.1 Data
    Metropolitan France housing price index comes from OECD [^4] The Organisation for Economic Co-operation and Development data and are from 1980 to 2021 (see OECD Stat, (2024) for details) in Base 2015. We renormalize in Base 2021 (i.e. in 2021, the housing price index is equal to $100$).
@@ 6.2 Definition of the climate transition
We consider four deterministic transition scenarios giving four deterministic carbon price trajectories. The scenarios used come from the NGFS [^5] Network for Greening the Financial System. simulations, whose descriptions are given by NGFS, (2022) as follows:
@@
    Net Zero 2050 is an ambitious scenario that limits global warming to $1.5^{\circ}C$ through stringent climate policies and innovation, reaching net zero $\mathrm{CO}_{2}$ [^6] Carbon dioxide. emissions around 2050. Some jurisdictions such as the United States of America, the European Union, and Japan reach net zero for all GHG by this point.
@@
For each scenario, the regulator decides the carbon price $P_{carbon,0}$ (in euro per ton of CO2-equivalent) at $t_{0}$ and the speed rate of the transition $\eta_{\delta}$. We summarize them in table 1 below.
@@
For the energy price, we consider electricity as the unique source of energy. Then, we assume a linear relation between the electricity and the carbon price inspired by Abrell et al., (2023), where a variation of the carbon price is linked withe the variation of the electricity by a the pass-through rate noted $k$  [^7] We can definitively consider other types of energy. Moreover, instead of working with a linear relation (with a pass-through rate) between the energy price and the carbon price, we could consider a more complex model such as NiGEM (the National Institute Global Econometric Model in Hantzsche et al., (2018)).. This means that $\mathfrak{f}^{elec}_{1}$ and $\mathfrak{f}^{elec}_{2}$ define in (4.32) are respectively $k$ and $P_{elec,0}-k\times P_{carbon,0}$. For France, we take the electricity price $P_{elec,0}=0.2161$ euro per Kilowatt-hour and $k=0.55$ (see Abrell et al., (2023)) ton per Kilowatt-hour. Its evolution is plotted in Figure 1(b).
@@ 6.3.2 Economic and housing pricing index (HPI) parameters
We keep the values of $\phi$, $\sigma$, $(\chi^{i})_{i\in\mathcal{I}}$, and $(\boldsymbol{\lambda}^{ji})_{i,j\in\mathcal{I}}$ already estimated. For the productivity process, we switched from a vector autoregressive process to an Ornstein-Uhlenbeck process. We therefore calibrate ${\mu}$, ${\varsigma}$, ${\Sigma}$, and ${\Gamma}$ as detailed in section 5.1.2 and we obtain the same results as in Sopgoui, (2024)[Section 4.3.1.].
@@ 6.4 Simulations and discussions
In the current simulations, since we are keeping the same data at the same frequency (annual), the main change is then to replace the Vector Autoregressive process by the O.-U. process. Therefore, the comments already made for (1) to (5) concerning the trends, the impact of the carbon price, the difference of scenarios, the relation between sectors, etc. do not change. We will focus here on the LGD and on the losses, with different type of collateral.
We will assume $D=500$. This may seem low but it is actually $500\times 500$ simulations. This is computationally-intensive. In fact, for each year $t$, we firstly simulate $500$ times $(\mathcal{Z}_{t},\mathcal{A}_{t})$. However, we are interested in one year risk measures  [^8] Recall for example that we are in 2025, we would like to compute the probability that a firm defaults between 1 January 2027 and 31 December 2027., therefore we also simulate $500$ times $(\mathcal{Z}_{t+1},\mathcal{A}_{t+1})$ given $(\mathcal{Z}_{t},\mathcal{A}_{t})$. Moreover, we considered 4 sectors and for each step, there are numerical integrals, matrices products, etc. to compute.
@@ 6.4.1 Impact of the carbon price on Loss Given Default
We compute here for $D=500$ simulations of the productivity processes $(\mathcal{Z},\mathcal{A})$, the loss given default of 4 loans with the same exposure but with 4 different financial assets collateral described in Table 3.
@@
We use the $D=500$ trajectories of the productivity processes $(\mathcal{Z},\mathcal{A})$ simulated above. We compute the loss given default of 4 loans with the same exposure but with the 4 buildings described in Table 5 as collateral.
@@
As we see in tables 6 and 3, all the comments made for a financial asset as collateral are valid for a building: the presence of a collateral reduces LGD, but this reduction is smaller when the climate transition scenario becomes tougher.
@@ 6.4.2 Expected and unexpected loss
We can remark that, for each $k=0,1,2$, firms $14k+$ 1, $24k+$ 2, $34k+$ 3, and $44k+$ 4 respectively belong to the Very High Emitting, High Emitting, Low Emitting, and Very Low Emitting groups. Moreover, we assume that
@@
We can finally say that introducing and increasing the carbon price in the economy will
  1. 1.
    increase bank charges (materialized by the level of provisions calculated from the expected loss $EL$) charged to clients or by the operating income of banks;
  2. 2.
    and reduce the solvency and profitability of banks (translated by the economic capital calculated from the unexpected loss $UL$).
For example, consider the Net Zero 2050 scenario, for a borrower belonging to a very high GHG-emitting sector and its collateral too, the bank will see its associated EL increases by 187% and its associated UL by 83%. Conversely, if the borrower is in a very low GHG emitting sector and its collateral too, the bank will obtain its associated EL increases by 11% and its UL unchanged.
@@ Conclusion
Following Bouveret et al., (2025), we developed here a framework to quantify the impacts of the carbon price on a credit portfolio (expected and unexpected) losses, when the obligor companies as well as their guarantees belong to an economy subject to the climate transition declined by carbon price. We start by describing a closed economy, driven by a productivity following a multidimensional Ornstein-Uhlenbeck and subject to a climate transition modeled through a dynamic and deterministic carbon price, by a dynamic stochastic multisectoral. Then, by using the discounted cash flow methodology with the cash flows, following a stochastic differential equation, depending on the productivity as well as the carbon price, we evaluate the obligor value that helps us later on to compute its probability of over-indebtedness. We then turn to the bank’s loss in the event of a borrower’s over-indebtedness and if its loan is collateralized. When that is the case, the potential loss of the bank is written as the difference between the debt amount (EAD) and the collateral liquidated. We finally distinguish two types of collateral: either a financial asset or a building, both belonging to the economy so affected by the productivity and the carbon price. This work opens the door to many extensions as a finer modeling of the real estate market, taking into account other types of guarantees, modeling the unsecured loans that we assumed constant, modelling the impact of the carbon price on the exposure. This work can be extended to include physical risk. In order to implement that, we could for example introduce the damage multiplier defined through the temperature, and which will affected the output of each sector. Therefore, for each sector $i\in\mathcal{I}$, we will have
$\exp{\left(-\frac{\gamma_{i}}{2}T_{t}^{2}\right)},$ (6.1)
where $\gamma_{i}\geq 0$ is the damage function coefficient and $T_{t}$ is the mean temperature of the earth relative to its pre-industrial level.